OpenStudy (anonymous):

Can someone explain how to find where the interval on a graph increases, decreases, and is constant?? PLEASE HELP

OpenStudy (anonymous):

do you have a picture, or a function ?

OpenStudy (anonymous):

a picture

OpenStudy (anonymous):

if you have a picture you do it with your eyeballs can you post a screenshot of the picture?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

when you answer questions about increasing or decreasing, answer in terms of the x values, not the y values in this case your curve is going down (decreasing) from \(-\infty\) to \(-2\)

OpenStudy (anonymous):

from \(x=-2\) to \(x=1\) the curve is going up (increasing)

OpenStudy (anonymous):

and it is constant (horizontal) after than, from \(1\) to \(\infty\)

OpenStudy (anonymous):

if you need to write the answer in interval notation, i am sure you can do that it is really more of a reading comprehension question, you have no equation to work with, just the picture hope it is more or less clear

OpenStudy (anonymous):

that helps a lot, thank you! can i ask two more questions about this graph?

OpenStudy (anonymous):

how to find range and relative minimum?

OpenStudy (anonymous):

nevermind i just found the range, so i just need relative min

OpenStudy (anonymous):

range, you look at the possible y values the little arrow pointing up on the left hand indicates that it goes up to \(\infty\) but the bottom most part of the curse is \(-4\) i think, making the range \[[-4,\infty)\]

OpenStudy (anonymous):

oh ok, is that what you got?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

k good relative min is \(-4\)

OpenStudy (anonymous):

in this case it is also the absolute min, but in general it doesn't have to be

OpenStudy (anonymous):

oh whoops. i got -4 for the relative min but the question i got wrong was "the number at which f has a relative min"

OpenStudy (anonymous):

oh, well i will stick with my answer, the relative min is -4 it occurs when x = -2, maybe that is what they were asking

OpenStudy (anonymous):

yep, that was it. i see it now. thank you so much!

OpenStudy (anonymous):

yw